# In-and-Out: Algorithmic Diffusion for Sampling Convex Bodies

### v2

@@ 3.4 Putting it together
For any given $\varepsilon\in(0,1)$ and set $\mathcal{K}\subset B_{D}(0)$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and any feasible start $x_{0}\in\mathcal{K}$ achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{d+D^{2}/h}{\varepsilon})$ iterations.
@@
$\mu_{1}^{X}(x)=\mathds{1}_{\mathcal{K}}(x)\cdot\int\frac{\exp(-\frac{1}{2h}\,\lVert y-x\rVert^{2})\exp(-\frac{1}{2h}\,\lVert y-x_{0}\rVert^{2})}{(2\pi h)^{d/2}\int_{\mathcal{K}}\exp(-\frac{1}{2h}\,\lVert y-x\rVert^{2})\,\mathrm{d}x}\,\mathrm{d}y\,.$
@@
Therefore, $M=\operatorname{ess\,sup}\frac{\mu_{1}^{X}}{\pi^{X}}\leq 2^{d/2}\exp(5h^{-1}D^{2})$. By Theorem 5 under (8), $\mathsf{In\text{-}and\text{-}Out}$ needs $\widetilde{\mathcal{O}}(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{d+D^{2}/h}{\varepsilon})$ iterations. ∎
@@
For any given $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset B_{D}(0)$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and an $M$-warm initial distribution achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qh^{-1}D^{2}\log\frac{1}{\varepsilon})$ iterations. If $\pi^{X}$ is isotropic, then $\mathsf{In\text{-}and\text{-}Out}$ only needs $\widetilde{\mathcal{O}}(qh^{-1}D\log\frac{d+d^{2}/h}{\varepsilon})$ iterations.
@@
For convex $\mathcal{K}$, it follows from Lemma 18 that $C_{\mathsf{LSI}}(\pi^{X})=\mathcal{O}(D^{2})$ and $C_{\mathsf{LSI}}(\pi^{X})=\mathcal{O}(D)$ for isotropic $\mathcal{K}$. The rest of the proof can be completed in a similar way. ∎
For $h=\tilde{\Theta}(d^{-2})$, $\mathsf{In\text{-}and\text{-}Out}$ requires $\widetilde{\mathcal{O}}(qd^{2}D^{2})$ iterations and in particular $\widetilde{\mathcal{O}}(qd^{2}D)$ iteration for isotropic uniform distributions. These results match those of $\mathsf{Speedy\ walk}$ [KLM06, LV17] (see Theorem 31).
@@ Acknowledgements.
Appendix A Related work
Sampling from constrained log-concave distributions is a fundamental task arising in many fields. Uniform sampling with convex constraints is its simplest manifestation, which was first studied as a core subroutine for a randomized volume-computation algorithm [DFK91]. Since then, this fundamental problem has been studied for over three decades [LS90, LS93, KLS97, LV03, LV06, BEL18, BDMP17]. We review these algorithms, grouping them under three categories — geometric random walks, structured samplers, and diffusion-type samplers. Below, $\mathcal{K}$ is convex.
Geometric random walk.
We discuss two geometric random walks – $\mathsf{Ball\ walk}$ [LS93, KLS97] and $\mathsf{Hit\text{-}and\text{-}Run}$ [Smi84, Lov99]. $\mathsf{Ball\ walk}$ is a simple metropolized random walk; it draws $y$ uniformly at random from a ball of radius $\delta$ centered at a current point $x$, and moves to $y$ if $y\in\mathcal{K}$ and stays at $x$ otherwise. In the literature, $\mathsf{Ball\ walk}$ actually refers to a composite algorithm consisting of [$\mathsf{Speedy\ walk}+$ rejection sampling], where $\mathsf{Speedy\ walk}$ records only the accepted steps of $\mathsf{Ball\ walk}$ (see §B for details). The step size $\delta$ should be set to $\mathcal{O}(d^{-1/2})$ to avoid stepping outside of $\mathcal{K}$. [KLS97] showed that $\mathsf{Ball\ walk}$ needs $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ membership queries to be $\varepsilon$-close to $\pi^{\mathcal{K}}$ in $\mathsf{TV}$, where $D$ is the diameter of $\mathcal{K}$, and the warmness parameter $M$ measures the closeness of the initial distribution to the target uniform distribution $\pi^{\mathcal{K}}$.
$\mathsf{Hit\text{-}and\text{-}Run}$ is another zeroth-order algorithm that needs no step size; it picks a uniform random line $\ell$ passing a current point, and move to a uniform random point on $\ell\cap\mathcal{K}$. [LV06] shows that, if we define the second moment as $R^{2}:=\mathbb{E}_{\pi^{\mathcal{K}}}[\lVert X-\mathbb{E}X\rVert^{2}]$, then $\mathsf{Hit\text{-}and\text{-}Run}$ requires $\mathcal{O}(d^{2}R^{2}\log\frac{M}{\varepsilon})$ queries. Notably, this algorithm has a poly-logarithmic dependence on $M$ as opposed to $\mathsf{Ball\ walk}$.
Both algorithm are affected by skewed shape of $\mathcal{K}$ (i.e., large $D$ or $R$), so these samplers are combined with pre-processing step called rounding. This procedure finds a linear transformation that makes the geometry of $\mathcal{K}$ less skewed and so more amenable to sampling. In literature, there exists a randomized algorithm [JLLV21] that rounds $\mathcal{K}$ and generates a good warm start (i.e., $M=\mathcal{O}(1)$), with $\mathsf{Ball\ walk}$ used as a core subroutine. We refer readers to [KZ24] for a streamlined proof, with $\mathsf{In\text{-}and\text{-}Out}$ used as a main sampler. This algorithm takes up $\widetilde{\mathcal{O}}(d^{3.5})$ queries in total, and in such position with the good warm start, $\mathsf{Ball\ walk}$ only needs $\widetilde{\mathcal{O}}(d^{2}\log\frac{1}{\varepsilon})$ queries to sample from $\pi^{\mathcal{K}}$.
Structured samplers.
The aforementioned samplers based on geometric random walks require only access to the membership oracle of the convex body without any additional structural assumptions. The alternate paradigm of geometry-aware sampling attempts to exploit the structure of convex constraints, with the aim of expediting the convergence of the resultant sampling schemes. One common assumption is to make available a self-concordant barrier function $\phi$ which has regularity on its high-order derivatives and blows up when approaching the boundary $\partial\mathcal{K}$. The Hessian of $\phi$ encodes the local geometry of the constraint, and the samplers often work directly with $\nabla^{2}\phi$.
The first canonical example of such a zeroth-order sampler is $\mathsf{Dikin\ walk}$ used when $\mathcal{K}$ is given by $m$ linear constraints [KN12]; it draws a uniform sample from an ellipsoid (characterized by $\nabla^{2}\phi$) of fixed radius around a current point, and is often combined with a Metropolis adjustment. [KN12] shows that $\mathsf{Dikin\ walk}$ mixes in $\mathcal{O}(md\log\frac{M}{\varepsilon})$ steps, although each iteration is slightly more expensive than one membership query. This algorithm requires no rounding, but still needs a good warm-start, which can be achieved by an annealing-type algorithm using $\widetilde{\mathcal{O}}(md)$ iterations of $\mathsf{Dikin\ walk}$ [KV24].
$\mathsf{Riemannian\ Hamiltonian\ Monte\ Carlo}$ is a structured sampler that exploits the first-order information of the potential (i.e., $\nabla\log(1/\pi)$) [GC11]; its proposal is given as the solution to the Hamilton’s ODE equation, followed by the Metropolis-filter. In the linear-constraint setting above, it requires $\mathcal{O}(md^{2/3}\log\frac{M}{\varepsilon})$ many iterations to achieve $\varepsilon$-close distance to $\pi^{\mathcal{K}}$ [LV18]. This sampler is further analyzed for practical ODE solvers [KLSV23] and for more sophisticated self-concordant barriers [GKV23].
Similarly, $\mathsf{Mirror\ Langevin}$ [ZPFP20, Jia21, AC21, LTVW22] is a class of algorithms which converts the constrained problem into an unconstrained one obtained by considering the pushforward of the constrained space by $\nabla\phi$. The algorithm can also be metropolized [SWW24]. The best known rate for this algorithm is $\widetilde{\mathcal{O}}(d\log\frac{1}{\varepsilon})$ under some strong assumptions on $\phi$.
Diffusion-type samplers.
Samplers based on discretizations of Itô diffusions, stochastic processes which rapidly mix to $\pi$ in continuous time, have long been used for sampling without constraints [BGHM95, DT12, Dal17, Che24]. While the underlying stochastic processes generalize easily to constrained settings, the discretization analysis relies crucially on the smoothness of the target distribution. This is clearly impossible to achieve in the constrained setting, so some techniques are required to circumvent this difficulty. These algorithms, however, generalize easily to the more general problem of sampling from distributions of the form $\tilde{\pi}^{X}\propto e^{-f}\mathds{1}_{\mathcal{K}}$, by incorporating first-order information from $f$.
The first approach for adapting diffusion-based samplers [BEL15, BEL18, Leh23] iterates a two-step procedure. First, a random step is taken, with $x_{k+1/2}\sim\mathcal{N}(x_{k},2hI_{d})$ for some appropriately chosen step $h$, [^4] A gradient step can be added in the more general case, for sampling from $\tilde{\pi}^{X}$. and then project it to $\mathcal{K}$, i.e., $x_{k+1}=\mathsf{proj}_{\mathcal{K}}(x_{k+1/2})$. The complexity is given in terms of queries to a projection oracle, each call to which can be implemented with a polynomial number of membership oracle queries; a total of $\tilde{\mathcal{O}}(\nicefrac{{d^{2}D^{3}}}{{\varepsilon^{4}}})$ queries are needed to be $\varepsilon$-close in $\mathcal{W}_{2}$ to $\pi^{X}$. Another approach, which uses an algorithmically designed “soft” penalty instead of a projection, was proposed in [GHZ22], and achieves a rate estimate of $\tilde{\mathcal{O}}(\nicefrac{{d}}{{\varepsilon^{10}}})$.
A second approach, suggested by [BDMP17], considers a different proximal scheme, which performs a “soft projection” onto $\mathcal{K}$, by taking steps like $\mathcal{N}((1-h\lambda^{-1})x_{k}+h\mathsf{proj}_{\mathcal{K}}(x_{k}),2hI_{d})$. It is called Moreau-Yosida regularized Langevin, named after an analogous regularization scheme for constrained optimization. This scheme also relies on access to a projection oracle for $\mathcal{K}$, and quantifies their query complexity accordingly. Their final rate estimate is $\tilde{\mathcal{O}}(\nicefrac{{d^{5}}}{{\varepsilon^{6}}})$ to be $\varepsilon$-close in $\mathsf{TV}$ distance to $\pi^{X}$.
Observing the prior work integrating diffusion-based sampling with convex constraints, the dependence on the key parameters $d,\varepsilon$, while polynomial, are many orders worse than the rates for zeroth-order samplers such as $\mathsf{Ball\ walk},\mathsf{Hit\text{-}and\text{-}Run}$. In contrast, our analysis not only recovers but in some sense surpasses the known rates for $\mathsf{Ball\ walk},\mathsf{Hit\text{-}and\text{-}Run}$, while harmonizing well with the continuous-time perspective of diffusions.
Proximal schemes for sampling.
The Gibbs sampling scheme used in this paper was inspired by the restricted Gaussian oracle introduced in [LST21] (in turn inspired by Gaussian Cooling [CV18]), which alternately iterates between a pure Gaussian step, and a “proximal” step (which we elaborate in our exposition). This scheme was given novel interpretations by [CCSW22], which showed that it interpolates the forward and backward heat flows, in the sense defined by [KP21]. The backward heat flow itself is intimately related to stochastic localization schemes, invented and popularized in [Eld13, Che21].
Appendix B $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$
@@
We restate the previously known guarantees for uniform sampling by $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$. Below, let $B_{r}(x)$ denote the $d$-dimensional ball of radius $r$ centered at $x$.
@@
$\mathsf{Ball\ walk}$ is particularly simple; draw a uniform random point from $B_{\delta}$ around the current point, and go there if the drawn point is inside of $\mathcal{K}$ and stay at the current point otherwise. Its stationary distribution can be easily seen to be $\pi\propto\mathds{1}_{\mathcal{K}}$, the uniform distribution over $\mathcal{K}$.
In the literature, there are two approaches to analyzing the convergence rate of this sampler: (i) a direct analysis via the $s$-conductance of $\mathsf{Ball\ walk}$ and (ii) an indirect approach which first passes through $\mathsf{Speedy\ walk}$.
@@
The following $\mathsf{TV}$-guarantee is obtained by lower bounding the $s$-conductance of $\mathsf{Ball\ walk}$, which requires a one-step coupling argument and the Cheeger inequality for $\pi$. We refer interested readers to [Vem05, §5].
@@
For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\pi_{t}$ be the distribution after $t$ steps of $\mathsf{Ball\ walk}$ with an $M$-warm initial distribution $\pi_{0}$. Then, $\mathsf{Ball\ walk}$ with step size $\delta=\Theta(\frac{\varepsilon}{M\sqrt{d}})$ achieves $\lVert\pi_{t}-\pi\rVert_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim d^{2}D^{2}\frac{M^{2}}{\varepsilon^{2}}\log\frac{M}{\varepsilon}$. If $\pi$ is isotropic, then $\mathsf{Ball\ walk}$ needs $\mathcal{O}(d^{2}\frac{M^{2}}{\varepsilon^{2}}\log d\log\frac{M}{\varepsilon})$ iterations.
The mixing time of $\mathsf{Ball\ walk}$ under this approach has a polynomial dependence on $1/\varepsilon$, rather than a polylogarithmic dependence.
@@
[KLS97] introduced $\mathsf{Speedy\ walk}$, which could be viewed as a version of $\mathsf{Ball\ walk}$ and converges to a speedy distribution (see Proposition 1), which is slightly biased from $\pi$. Then, $\mathsf{Speedy\ walk}$ is used together with another algorithmic component (rejection sampling) [KLS97, Algorithm 4.15] that converts the speedy distribution to the uniform distribution. In the literature, $\mathsf{Ball\ walk}$ often refers to ‘$\mathsf{Speedy\ walk}$ combined with the conversion step’, rather than a direct implementation of Algorithm 2. Strictly speaking, a mixing guarantee of this combined algorithm should not be referred to as a provable guarantee of $\mathsf{Ball\ walk}$.
@@
As opposed to $\mathsf{Ball\ walk}$, $\mathsf{Speedy\ walk}$ always takes some step at each iteration. However, the problem of sampling from $x_{i}\sim\textnormal{Unif}\mathopen{}\mathclose{{}\left(\mathcal{K}\cap B_{\delta}(x_{i-1})}\right)$ in Line 3 is not straightforward. This step admits the following implementation based on rejection sampling, via a procedure denoted by ()($\ast$)( ∗ ):
@@
Each actual step (indexed by $i$) in $\mathsf{Speedy\ walk}$ is called a proper step, and rejected steps during ()($\ast$)( ∗ ) are called improper steps. For example, if $x_{1},x_{1},x_{2},x_{3},x_{3},x_{3},x_{4},\dotsc$ are the positions produced by $\mathsf{Ball\ walk}$, then only proper steps $x_{1},x_{2},x_{3},x_{4},\dotsc$ are recorded by $\mathsf{Speedy\ walk}$.
To describe the theoretical guarantees of $\mathsf{Speedy\ walk}$, we define the local conductance $\ell(x)$ at $x\in\mathcal{K}$, which measures the success probability of the rejection sampling scheme in ()($\ast$)( ∗ ):
@@
The stationary distribution $\nu$ of $\mathsf{Speedy\ walk}$ has density
[…]
@@
The best known result for $\mathsf{Speedy\ walk}$’s mixing is due to [KLM06] devising the blocking conductance and using the log-Cheeger inequality. When $\nu$ is isotropic (i.e., it has covariance proportional to the identity matrix), [LV17] improves the mixing bound via the log-Cheeger constant.
@@
For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\nu_{t}$ be the distribution after $t$ proper steps of $\mathsf{Speedy\ walk}$ started at any feasible point $x_{0}\in\mathcal{K}$. Then, $\mathsf{Speedy\ walk}$ with step size $\delta=\Theta(d^{-1/2})$ achieves $\lVert\nu_{t}-\nu\rVert_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim(D^{2}+\log(D\sqrt{d}))\,d^{2}\log\frac{1}{\varepsilon}$. From an $M$-warm start, the expected number of improper steps during $t$ iterations is $\widetilde{\mathcal{O}}(tM)$. When $\nu$ is isotropic, $\mathsf{Speedy\ walk}$ needs $\mathcal{O}(d^{2}D\log\frac{1}{\varepsilon}\log\log D)$ proper steps to achieve $\varepsilon$-$\mathsf{TV}$ distance to $\nu$.
@@
$\mathcal{A}$: Call $\mathsf{Speedy\ walk}$ to obtain a sample $X\sim\nu_{t}$ until $\frac{2d}{2d-1}\,X\in\mathcal{K}$. If so, return $\bar{X}=\frac{2d}{2d-1}\,X$.
@@
Combining the previous two results, we conclude that the total expected number of membership queries to obtain a sample $\varepsilon$-close to $\pi$ in $\mathsf{TV}$ is $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$, which now has a poly-logarithmic dependence on $1/\varepsilon$.
Remark 3 (Backward heat flow analysis of $\mathsf{Speedy\ walk}$).
Consider a Gaussian version of $\mathsf{Speedy\ walk}$, whose one-step corresponds to $x_{i+1}\sim\mathcal{N}(x_{i},hI_{d})|_{\mathcal{K}}$, and this transition kernel exactly matches integrating ($\mathsf{BH}$) for time $h$. Thus, $\nu Q^{\pi^{X},h}_{h}=\nu$ due to the stationarity of $\nu$ under $\mathsf{Speedy\ walk}$, where $Q^{\pi^{X},h}_{h}$ is the transition kernel defined by the backward heat flow for time $h$ that reverses $\pi^{X}*\mathcal{N}(0,hI_{d})$ to $\pi^{X}$. Hence, if we can control the LSI/PI constants of $\nu$ along the backward heat-flow’s trajectory, then we could directly analyze $\mathsf{Speedy\ walk}$ by emulating computations in Lemma 21.
Appendix C Functional inequalities
We provide full details on functional inequalities omitted in §3.1. We use $\mu$ and $\mu_{\mathsf{LC}}$ to denote a probability measure and log-concave probability measure over $\mathbb{R}^{d}$, respectively.

### v3

@@ 3.4 Putting it together
For any given $\varepsilon\in(0,1)$ and set $\mathcal{K}\subset B_{D}(0)$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and any feasible start $x_{0}\in\mathcal{K}$ achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{O}(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{d+D^{2}/h}{\varepsilon})$ iterations.
@@
$\mu_{1}^{X}(x)=\mathds{1}_{\mathcal{K}}(x)\int\frac{\exp(-\frac{1}{2h}\,\lvert y-x\rvert^{2})\exp(-\frac{1}{2h}\,\lvert y-x_{0}\rvert^{2})}{(2\pi h)^{d/2}\int_{\mathcal{K}}\exp(-\frac{1}{2h}\,\lvert y-x\rvert^{2})\,\mathrm{d}x}\,\mathrm{d}y\,.$
@@
Therefore, $M=\operatorname{ess\,sup}\frac{\mu_{1}^{X}}{\pi^{X}}\leq 2^{d/2}\exp(5h^{-1}D^{2})$. By Theorem 5 under (8), $\mathsf{In\text{-}and\text{-}Out}$ needs $\widetilde{O}(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{d+D^{2}/h}{\varepsilon})$ iterations. ∎
@@
For any given $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset B_{D}(0)$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and a feasible start $x_{0}\in\mathcal{K}$ achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{O}(qh^{-1}D^{2}\log\frac{1}{\varepsilon})$ iterations. If $\pi^{X}$ is isotropic, then $\mathsf{In\text{-}and\text{-}Out}$ only needs $\widetilde{O}(qh^{-1}D\log\frac{d+d^{2}/h}{\varepsilon})$ iterations.
@@
For convex $\mathcal{K}$, it follows from Lemma 17 that $C_{\mathsf{LSI}}(\pi^{X})=O(D^{2})$ and $C_{\mathsf{LSI}}(\pi^{X})=O(D)$ for isotropic $\mathcal{K}$. The rest of the proof can be completed in a similar way. ∎
For $h=\tilde{\Theta}(d^{-2})$, $\mathsf{In\text{-}and\text{-}Out}$ requires $\widetilde{O}(qd^{2}D^{2})$ iterations and in particular $\widetilde{O}(qd^{2}D)$ iteration for isotropic uniform distributions. These bounds match those of the $\mathsf{Speedy\ walk}$ [KLM06, LV24] (see Theorem 27), albeit with stronger guarantees on the output distribution.
@@ Acknowledgements.
Appendix A $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$
@@
We restate the previously known guarantees for uniform sampling by the $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$. Below, let $B_{r}(x)$ denote the $d$-dimensional ball of radius $r$ centered at $x$.
@@
The $\mathsf{Ball\ walk}$ is particularly simple; draw a uniform random point from $B_{\delta}$ around the current point, and go there if the drawn point is inside of $\mathcal{K}$ and stay at the current point otherwise. Its stationary distribution can be easily seen to be $\pi\propto\mathds{1}_{\mathcal{K}}$, the uniform distribution over $\mathcal{K}$.
In the literature, there are two approaches to analyzing the convergence rate of this sampler: (i) a direct analysis via the $s$-conductance of the $\mathsf{Ball\ walk}$ and (ii) an indirect approach which first passes through the $\mathsf{Speedy\ walk}$.
@@
The following $\mathsf{TV}$-guarantee is obtained by lower bounding the $s$-conductance of the $\mathsf{Ball\ walk}$, which requires a one-step coupling argument and the Cheeger inequality for $\pi$. We refer interested readers to [Vem05, §5].
@@
For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\pi_{t}$ be the distribution after $t$ steps of the $\mathsf{Ball\ walk}$ with an $M$-warm initial distribution $\pi_{0}$. Then, the $\mathsf{Ball\ walk}$ with step size $\delta=\Theta(\frac{\varepsilon}{M\sqrt{d}})$ achieves $\lVert\pi_{t}-\pi\rVert_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim d^{2}D^{2}\frac{M^{2}}{\varepsilon^{2}}\log\frac{M}{\varepsilon}$. If $\pi$ is isotropic, then the $\mathsf{Ball\ walk}$ needs $O(d^{2}\frac{M^{2}}{\varepsilon^{2}}\log d\log\frac{M}{\varepsilon})$ iterations.
The mixing time of the $\mathsf{Ball\ walk}$ under this approach has a polynomial dependence on $1/\varepsilon$, rather than a polylogarithmic dependence.
@@
[KLS97] introduced the $\mathsf{Speedy\ walk}$, which could be viewed as a version of the $\mathsf{Ball\ walk}$ and converges to a speedy distribution (see Proposition 2), which is slightly biased from $\pi$. Then, the $\mathsf{Speedy\ walk}$ is used together with another algorithmic component (rejection sampling) [KLS97, Algorithm 4.15] that converts the speedy distribution to the uniform distribution. In the literature, the $\mathsf{Ball\ walk}$ often refers to $\mathsf{Speedy\ walk}$ combined with the conversion step’, rather than a direct implementation of Algorithm 2. Strictly speaking, a mixing guarantee of this combined algorithm should not be referred to as a provable guarantee of the $\mathsf{Ball\ walk}$.
@@
As opposed to the $\mathsf{Ball\ walk}$, the $\mathsf{Speedy\ walk}$ always takes some step at each iteration. However, the problem of sampling from $x_{i}\sim\textnormal{Unif}\mathopen{}\mathclose{{\left(\mathcal{K}\cap B_{\delta}(x_{i-1})}}\right)$ in Line 3 is not straightforward. This step admits the following implementation based on rejection sampling, via a procedure denoted by $(\ast)$:
@@
Each actual step (indexed by $i$) in the $\mathsf{Speedy\ walk}$ is called a proper step, and rejected steps during $(\ast)$ are called improper steps. For example, if $x_{1},x_{1},x_{2},x_{3},x_{3},x_{3},x_{4},\dotsc$ are the positions produced by the $\mathsf{Ball\ walk}$, then only proper steps $x_{1},x_{2},x_{3},x_{4},\dotsc$ are recorded by the $\mathsf{Speedy\ walk}$.
To describe the theoretical guarantees of the $\mathsf{Speedy\ walk}$, we define the local conductance $\ell(x)$ at $x\in\mathcal{K}$, which measures the success probability of the rejection sampling scheme in $(\ast)$:
@@
The stationary distribution $\nu$ of the $\mathsf{Speedy\ walk}$ has density
$\nu(x)=\frac{\ell(x)\,\mathds{1}_{\mathcal{K}}(x)}{\int_{\mathcal{K}}\ell(x)\,\mathrm{d}x}\,.$
@@
The best known result for $\mathsf{Speedy\ walk}$’s mixing is due to [KLM06] devising the blocking conductance and using the log-Cheeger inequality. When $\nu$ is isotropic (i.e., it has covariance proportional to the identity matrix), [LV24] improves the mixing bound via the log-Cheeger constant.
@@
For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\nu_{t}$ be the distribution after $t$ proper steps of the $\mathsf{Speedy\ walk}$ started at any feasible point $x_{0}\in\mathcal{K}$. Then, the $\mathsf{Speedy\ walk}$ with step size $\delta=\Theta(d^{-1/2})$ achieves $\lVert\nu_{t}-\nu\rVert_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim(D^{2}+\log(D\sqrt{d}))\,d^{2}\log\frac{1}{\varepsilon}$. From an $M$-warm start, the expected number of improper steps during $t$ iterations is $\widetilde{O}(tM)$. When $\nu$ is isotropic, the $\mathsf{Speedy\ walk}$ needs $\mathcal{O}(d^{2}D\log\frac{1}{\varepsilon}\log\log D)$ proper steps to achieve $\varepsilon$-$\mathsf{TV}$ distance to $\nu$.
@@
$\mathcal{A}$: Call the $\mathsf{Speedy\ walk}$ to obtain a sample $X\sim\nu_{t}$ until $\frac{2d}{2d-1}\,X\in\mathcal{K}$. If so, return $\bar{X}=\frac{2d}{2d-1}\,X$.
@@
Combining the previous two results, we conclude that the total expected number of membership queries to obtain a sample $\varepsilon$-close to $\pi$ in $\mathsf{TV}$ is $\widetilde{O}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$, which now has a poly-logarithmic dependence on $1/\varepsilon$.
Appendix B Functional inequalities
We provide full details on functional inequalities omitted in §3.1. We use $\mu$ and $\mu_{\mathsf{LC}}$ to denote a probability measure and logconcave probability measure over $\mathbb{R}^{d}$, respectively.